# Capacity-Aware Inference: Mitigating the Straggler Effect in Mixture of Experts

### v1

@@ Abstract
The Mixture of Experts (MoE) is an effective architecture for scaling large language models by leveraging sparse expert activation, optimizing the trade-off between performance and efficiency. However, under expert parallelism, MoE suffers from inference inefficiencies due to imbalanced token-to-expert assignment, where some experts are overloaded while others remain underutilized. This imbalance leads to poor resource utilization and increased latency, as the most burdened expert dictates the overall delay, a phenomenon we define as the Straggler Effect. To mitigate this, we propose Capacity-Aware Inference, including two key techniques: (1) Capacity-Aware Token Drop, which discards overloaded tokens to regulate the maximum latency of MoE, and (2) Capacity-Aware Token Reroute, which reallocates overflowed tokens to underutilized experts, balancing the token distribution. These techniques collectively optimize both high-load and low-load expert utilization, leading to a more efficient MoE inference pipeline. Extensive experiments demonstrate the effectiveness of our methods, showing significant improvements in inference efficiency, e.g., 0.2% average performance increase and a 1.94$\times$ inference speedup on Mixtral-8$\times$7B-Instruct.
Capacity-Aware Inference: Mitigating the Straggler Effect in
Mixture of Experts
Shwai He1 Weilin Cai2 Jiayi Huang2 Ang Li1 1University of Maryland, College Park 2The Hong Kong University of Science and Technology (Guangzhou) shwaihe@umd.edu, angliece@umd.edu
@@ 1 Introduction
In recent years, the rapid evolution of Large Language Models (LLMs) OpenAI (2024); Team (2024a); DeepSeek-AI et al. (2024b) has driven a wave of innovations, continuously expanding the frontiers of AI research and applications. Among the model architectural innovations, the Mixture of Experts (MoE) framework has emerged as a pivotal technique for optimizing the cost-performance trade-off in LLMs. Specifically, MoE Shazeer et al. (2017a); Fedus et al. (2022) enhances scalability by integrating multiple experts while activating only a subset per input. This selective activation substantially improves model performance without a corresponding increase in computational cost, effectively balancing efficiency and performance.
Despite the success of MoE, a key efficiency challenge lies in the imbalanced token-to-expert distribution, which results in some experts being overloaded while others remain underutilized Lepikhin et al. (2021); Zoph et al. (2022). In the context of expert parallelism, low-load experts complete their computations faster but must wait for high-load experts to finish, leading to inefficient resource utilization and increased latency. As illustrated in Figure 1, this phenomenon is defined as the Straggler Effect, where the most burdened expert dictates the overall latency of the MoE layer.
@@
Figure 1: Illustration of the Straggler Effect in MoE Inference, where the most burdened expert (i.e., Straggler Expert) dictates the overall latency.
@@
While auxiliary balance losses have been incorporated into the training process to alleviate imbalance Shazeer et al. (2017a); Fedus et al. (2022); DeepSeek-AI et al. (2024b), these techniques remain ineffective in mitigating imbalance during inference. Specifically, as shown in Figure 1, our findings reveal a highly uneven token distribution among experts, with the highest-load expert handling more than seven times the expected average load. Moreover, managing such an imbalance during inference often incurs additional resource costs. For instance, DeepSeek-V3 addresses this issue by duplicating high-load experts and deploying them redundantly DeepSeek-AI et al. (2024b). This motivates us to explore efficient token-to-expert assignment through two key questions: (1) How to prevent extreme overloading in high-load experts? and (2) How can we utilize the available capacity of low-load experts?
We address these challenges through Capacity-Aware Inference. For high-load experts, we introduce Capacity-Aware Token Drop, a technique that removes excessive tokens from overloaded experts. On the one hand, eliminating these excess tokens alleviates extreme load imbalances, significantly enhancing efficiency. On the other hand, dropping these tokens minimally impacts model performance, as they constitute only a small fraction of the total tokens. For low-load experts, we extend Token Drop with Capacity-Aware Token Reroute, which leverages the available capacity of underutilized experts to process overflowed tokens. Token Reroute further enhances the performance of Token Drop within the capacity-constrained inference framework. Extensive experimental results validate the effectiveness of our proposed techniques, demonstrating significant improvements in both efficiency and performance, e.g., 0.2% average performance increase and a 1.94$\times$ inference speedup in Mixtral-8$\times$7B-Instruct. In short, our contributions are in three folds:
@@
  • Toward token imbalance, we propose Token Drop and Token Reroute to enhance the balanced utilization of experts.
  • Experimental results demonstrate the effectiveness of Token Drop and Token Reroute, achieving significant improvements in expert utilization and inference efficiency.
@@ 2 Related Works
The Mixture of Experts (MoE) is a kind of neural network architecture with an extended set of parameters (referred to as “experts”) controlled by a router, which is first introduced in the context of conditional computation (Jacobs et al., 1991; Jordan and Jacobs, 1994). The potential of sparse activation in MoE is subsequently exploited by Shazeer et al. (2017b) for efficient training and inference on pretrained models with special designs, opening the door for MoE in various vision (Riquelme et al., 2021) and language (Lepikhin et al., 2020; Du et al., 2022; Fedus et al., 2022) scenarios. Attributed to its exceptional efficiency, MoE has been adopted as a foundational framework in the designs of large language models (Jiang et al., 2024; Dai et al., 2024; Xue et al., 2024a; Zhu et al., 2024; Team, 2024b), achieving superior scaling laws at low computational costs. Despite these advancements, MoE still faces efficiency challenges in both training and inference Cai et al. (2024), and our work specifically focuses on enhancing inference-time efficiency.
@@
The imbalance in token-to-expert assignments Zhou et al. (2022); Chen et al. (2022) poses a significant challenge to the deployment of Mixture of Experts (MoE). This imbalance leads to inefficiencies in memory, computation, and communication (He et al., 2023; Song et al., 2023; Xue et al., 2024b), making it a critical bottleneck for MoE scalability and deployment. To mitigate this issue, an auxiliary balance loss Shazeer et al. (2017a) is incorporated into the training process to encourage more uniform token distribution across experts. Additionally, various training strategies have been introduced to further balance token assignments: Switch-Transformer Fedus et al. (2022) and DeepSeek-V2 DeepSeek-AI et al. (2024a) implement Token Drop to alleviate expert overload, while DeepSeek-V3 DeepSeek-AI et al. (2024b) introduces an additional sequence-level auxiliary loss to prevent severe token imbalance.
However, these techniques primarily focus on training and fail to ensure balanced token assignments during inference. Instead, addressing token imbalance at inference often incurs additional resource costs. For example, DeepSeek-V3 DeepSeek-AI et al. (2024b) mitigates this issue by duplicating high-load experts and deploying them redundantly. In contrast, our approach effectively balances token assignments without introducing additional computational overhead.
@@ 3.1 Extremely Imbalanced Expert Utilization
While experts in MoE can be deployed in parallel, imbalanced token-to-expert assignments lead to varying levels of expert utilization and introduce potential latency. Despite the incorporation of balancing techniques during training, the load imbalance persists during inference. To further investigate this issue, we conduct preliminary experiments to analyze expert-specific utilization patterns and assess the impact of imbalance on practical latency.
@@
Figure 2 illustrates the normalized peak tokens load for each expert to accommodate all tokens within a single layer of OLMoE, where some experts receive an excessively large number of tokens (e.g., more than seven times of average load), which resulting in significant latency. A detailed layer-by-layer analysis is provided in Appendix B.
@@ 3.2 Motivation – the Straggler Effect
where $N_{i}$ represents the number of tokens assigned to the $i$-th expert, with the total token allocation satisfying $\sum\nolimits_{i=1}^{n}N_{i}=tk$. According to Eq. 4, the latency follows the Straggler Effect: the most burdened expert dictates the overall latency of the MoE layer. In the worst case, all tokens are assigned to the same group of experts, underutilizing the parallel processing capability of MoE. Conversely, distributing tokens evenly across experts maximizes computational efficiency and fully leverages the parallelism of multiple experts. With the bounds of the ideal and worst cases, the range of the highest load is given by:
@@
Given that the imbalance stems from excessively high- and low-load experts, we address this issue by exploring the following questions: (1) Are there redundant tokens assigned to High-Load Experts? That is, can the MoE model maintain its performance without these excess tokens? (2) Do Low-Load Experts matter? Should these experts be removed, or should their utilization be increased to sustain or even enhance overall efficacy?
@@ 4 Methodology
(a) Capacity-Aware Token Drop
(b) Capacity-Aware Token Reroute
Figure 3: Illustrations of (a) Capacity-Aware Token Drop and (b) Capacity-Aware Token Reroute. Token Drop mitigates overload by discarding excess tokens from overloaded experts, while Token Reroute redistributes these tokens to underutilized experts, enhancing load balancing.
@@
where $\gamma$ is typically much smaller than $\frac{n}{k}$. This constraint ensures that no expert exceeds the specified capacity limit, effectively mitigating severe load imbalances and reducing latency.
@@
where $s_{ij}$ denotes the importance score of the mapping from the $i$-th token to the $j$-th expert. With this score, each overflowed expert selectively discards those with lower scores. For the $j$-th expert:
$K_{j}=N_{j}-C,$ (9)
$\tau_{j}=\text{KthValue}(\mathcal{S}[:,j],K_{j}),$ (10)
where $K$ denotes the number of overflowed tokens, $\tau$ represents the $K$-th smallest value in $\mathcal{S}[:,j]$, serving as a threshold to filter out excess tokens:
$T_{j}\leftarrow\{t\mid\mathcal{S}[t,j]\leq\tau_{j}\}$ (11)
$\mathcal{S}[T_{j},j]\leftarrow 0,$ (12)
where $T_{j}$ denotes the indices of the rejected tokens from the $j$-th Expert. We mask the scores of the rejected tokens and prevent them from being sent to their corresponding overflowed experts.
Regarding the specific scoring function, we explore multiple efficient metrics and summarize them as follows:
Order: Discarding later tokens once earlier tokens have filled the expert capacity. This strategy was first introduced in Switch-Transformer Fedus et al. (2022) during training, and we extend it to the inference phase.
@@
Among these metrics, “Order” and “Reverse Order” are unstable, as shuffling sequences within a batch may result in different tokens being dropped Hayes et al. (2024). “Random” assumes all tokens have an equal probability of being dropped. In contrast, “Score” is stable, unaffected by sequence order within a batch.
Algorithm 1 Capacity-Aware Token Reroute
1:Expert Capacity $C$, Rounds $R$
2:Updated Scores $\mathcal{S}$
3:for $r=1,\dots,R$ do
4:     $\boldsymbol{M}\leftarrow\text{Top-$k$}(\mathcal{S},k)$
5:     $\triangleright$ Indicator for top-$k$ mappings
6:     $\mathcal{S}_{r}\leftarrow\mathcal{S}\odot\boldsymbol{M}$
7:     $\triangleright$ Mask out the non-selected
8:     for $j=1,\dots,n$ do
9:         $N_{j}\leftarrow||\mathcal{S}_{r}[:,j]||_{0}$
10:         $\triangleright$ Compute the load
11:         $K_{j}\leftarrow N_{j}-C$
12:         $\triangleright$ Number of excess tokens
13:         if $k>0$ then
14:              $\tau_{j}\leftarrow\text{KthValue}(\mathcal{S}_{r}[:,j],K_{j})$
15:              $\triangleright$ Threshold for filtering
16:              $T_{j}\leftarrow\{t\mid\mathcal{S}_{r}[t,j]\leq\tau_{j}\}$
17:              $\triangleright$ Indices of retained tokens
18:              $\mathcal{S}[T_{j},j]\leftarrow 0$
19:              $\triangleright$ Discard overflowed assignments
20:         end if
21:     end for
22:end for
Token Reroute Enhances the Utilization of Low-load Experts
Token Drop exclusively targets overloaded experts by discarding overflowed tokens that exceed expert capacity but does not address the underutilization of low-load experts. Next, we introduce Token Reroute to ensure a more balanced token-to-expert allocation. For tokens rejected by overflowed experts, we reselect underutilized experts based on the updated importance scores following Token Drop. As shown in Equation 12, masking the mappings to overflowed experts prevents these tokens from being reassigned to them, thereby encouraging their redistribution to other available experts. As illustrated in Algorithm 1, the process of masking and recomputing can be repeated iteratively, gradually achieving a more balanced token-to-expert assignment. Notably, the extra cost of Token Routing is minimal, as it occurs before tokens are assigned to each expert and operates solely on the lightweight importance matrix. Additionally, overflowed experts and tokens that have already selected $k$ experts are excluded from subsequent iterations, further minimizing the additional computational cost of the Top-$k$ operation.

### v3

@@ Abstract
The Mixture of Experts (MoE) is an effective architecture for scaling large language models by leveraging sparse expert activation to balance performance and efficiency. However, under expert parallelism, MoE suffers from inference inefficiencies due to imbalanced token-to-expert assignment, where underloaded experts complete computations early but must wait for overloaded experts, leading to global delays. We define this phenomenon as the Straggler Effect, as the most burdened experts dictate the overall inference latency. To address this, we first propose Capacity-Aware Token Drop, which enforces expert capacity limits by discarding excess tokens from overloaded experts, effectively reducing load imbalance with minimal performance impact (e.g., $30\%$ speedup with only $0.9\%$ degradation on OLMoE). Next, given the presence of low-load experts remaining well below the capacity threshold, we introduce Capacity-Aware Expanded Drop, which allows tokens to include additional local experts in their candidate set before enforcing strict local capacity constraints, thereby improving load balance and enhancing the utilization of underused experts. Extensive experiments on both language and multimodal MoE models demonstrate the effectiveness of our approach, yielding substantial gains in expert utilization, model performance, and inference efficiency, e.g., applying Expanded Drop to Mixtral-8$\times$7B-Instruct yields a 0.2% average performance improvement and a 1.85$\times$ inference speedup.
@@ 1 Introduction
Figure 1: Illustration of the Straggler Effect in MoE Inference, where the most burdened experts dictate the overall latency.
In recent years, the rapid evolution of Large Language Models (LLMs) [32, 38, 11] has driven a wave of innovations, continuously expanding the frontiers of AI research and applications. Among the model architectural innovations, the Mixture of Experts (MoE) framework has emerged as a pivotal technique for optimizing the cost-performance trade-off in LLMs. Specifically, MoE [35, 14] enhances scalability by integrating multiple experts while activating only a subset per input. This selective activation substantially improves model performance without a corresponding increase in computational cost, effectively balancing efficiency and performance.
Despite the success of MoE, a key efficiency challenge lies in the imbalanced token-to-expert distribution, which results in some experts being overloaded while others remain underutilized [26, 46]. In distributed GPU settings, experts are typically sharded across multiple devices, with each GPU responsible for a subset of the experts. Under expert parallelism, low-load experts complete their computations earlier but must wait for overloaded experts to finish, as synchronization barriers are required before proceeding to the next stage. This expert-level straggler effect further propagates to device-level delays, where GPUs hosting lighter expert workloads are stalled by GPUs hosting heavier workloads, leading to inefficient resource utilization and increased end-to-end latency during inference. As illustrated in Figure 1, this phenomenon is referred to as the Straggler Effect, where the heavily loaded experts determine the overall latency of imbalanced MoE inference.
@@
While auxiliary balance losses have been incorporated into the training process to alleviate imbalance [35, 14, 11], these techniques remain ineffective in mitigating imbalance during inference. Specifically, as shown in Figure 1, our findings reveal a highly uneven token distribution among experts, with the highest-load expert handling more than seven times the expected average load. Moreover, managing such imbalance during inference often incurs additional resource overhead. For example, DeepSeek-V3 mitigates this issue by duplicating high-load experts and deploying them redundantly across devices [11]. This motivates us to explore efficient token-to-expert assignment by addressing the key question: How can we prevent extreme overloading of heavily utilized experts?
We propose Capacity-Aware Inference to address this challenge. Specifically, for high-load experts, we introduce Capacity-Aware Token Drop, which imposes a maximum capacity constraint and discards excess tokens from overloaded experts. This approach alleviates severe load imbalance and significantly improves efficiency, while maintaining model performance since the dropped tokens represent only a small fraction of the total workload, e.g., OLMoE achieves a 30% speedup in MoE layers with just a 0.9% performance degradation. After removing excess tokens from overloaded experts, we observe that some low-load experts remain significantly underutilized relative to the predefined capacity constraints, yet must still wait for other experts to complete their computations. This leads us to a second key question: How can we effectively leverage the available capacity of underutilized experts?
For low-load experts, we extend Token Drop with Capacity-Aware Expanded Drop, which further utilizes the available capacity of underutilized experts to handle overflow tokens from high-load experts. Specifically, under expert parallelism across multiple GPUs, Expanded Drop allows each token to consider additional candidate experts on the same device while still enforcing strict local capacity constraints. This expanded selection improves the utilization of low-load experts and enhances the representational capacity of MoE models under capacity-constrained scenarios.
Extensive experimental results validate the effectiveness of our proposed techniques, demonstrating significant improvements in both efficiency and performance, e.g., applying Expanded Drop to Mixtral-8$\times$7B-Instruct yields a 0.2% average performance improvement and a 1.85$\times$ inference speedup. Moreover, in multimodal models, we identify redundancy among image tokens and show that applying aggressive capacity constraints (e.g., setting the maximum to half of the average expert load) can still maintain performance. In short, our contributions are in four folds:
@@
  • Toward the high-load experts, we propose Capacity-Aware Token Drop, which enforces capacity constraints by discarding excess tokens assigned to overloaded experts, thereby mitigating extreme load imbalance.
  • To better utilize underloaded experts, we introduce Capacity-Aware Expanded Drop, which expands the candidate expert set to include additional local experts, further improving load balance and model performance.
  • Extensive experiments on both language and multimodal models validate the effectiveness of our approach, demonstrating substantial improvements in inference efficiency with comparable performance.
@@ 2 Related Works
The Mixture of Experts (MoE) is a kind of neural network architecture with an extended set of parameters (referred to as “experts”) controlled by a router, which is first introduced in the context of conditional computation [21, 23]. The potential of sparse activation in MoE is subsequently exploited by [34] for efficient training and inference on pretrained models with special designs, opening the door for MoE in various vision [33] and language [25, 13, 14] scenarios. Attributed to its exceptional efficiency, MoE has been adopted as a foundational framework in the designs of large language models [22, 9, 41, 45, 39], achieving superior scaling laws at low computational costs. Despite these advancements, MoE still faces efficiency challenges in both training and inference [3], and our work specifically focuses on enhancing inference-time efficiency.
@@
The imbalance in token-to-expert assignments [44, 5] poses a significant challenge to the deployment of MoE. This imbalance leads to inefficiencies in computation, communication, and memory [18, 37, 42], making it a critical bottleneck for MoE scalability and deployment. To mitigate this issue, an auxiliary balance loss [35] is incorporated into the training process to encourage more uniform token distribution across experts. Additionally, various training strategies have been introduced to further balance token assignments: Switch-Transformer [14] and DeepSeek-V2 [10] implement Token Drop to alleviate expert overload, while DeepSeek-V3 [11] introduces an additional sequence-level auxiliary loss to prevent severe token imbalance.
However, these techniques primarily focus on training and fail to ensure balanced token assignments during inference. Instead, addressing token imbalance at inference often incurs additional resource costs. For example, DeepSeek-V3 [11] mitigates this issue by duplicating high-load experts and deploying them redundantly. In contrast, our approach effectively balances token assignments without introducing additional computational overhead.
@@ 3.1 Extremely Imbalanced Expert Utilization
While experts in MoE models are often deployed in parallel across distributed GPUs, imbalanced token-to-expert assignments lead to varying levels of expert utilization and potential latency. Despite the incorporation of balancing techniques during training, the load imbalance persists during inference. To further investigate this issue, we conduct preliminary experiments to analyze expert-specific utilization patterns and assess the impact of imbalance on practical latency.
@@
Figure 2 illustrates the normalized peak tokens load for each expert to accommodate all tokens within a single layer of OLMoE, where some experts receive an excessively large number of tokens (e.g., more than seven times the average number of tokens), leading to severe load imbalance and, consequently, significant latency. A detailed layer-by-layer analysis is provided in Appendix E.
@@ 3.2 Motivation – the Straggler Effect
where $N_{i}$ represents the number of tokens assigned to the $i$-th expert, with the total token allocation satisfying $\sum\nolimits_{i=1}^{n}N_{i}=tk$. According to Eq. 4, the latency follows the Straggler Effect: the most burdened experts dictate the overall latency of the MoE layer. In the worst case, all tokens are assigned to the same group of experts, underutilizing the parallel processing capability of MoE. Conversely, distributing tokens evenly across experts maximizes computational efficiency and fully leverages the parallelism of multiple experts. With the bounds of the ideal and worst cases, the range of the highest load is given by:
@@
Given that the imbalance stems from excessively high- and low-load experts, we address this issue by exploring the following questions: (1) For high-load experts, are there redundant tokens that can be dropped without causing significant performance degradation? (2) For low-load experts that must wait for high-load experts to complete forward passes, is there an opportunity to enhance their utilization and improve performance without incurring substantial additional cost?
@@ 4 Methodology
where $\gamma$ is typically much smaller than $\frac{n}{k}$. This constraint ensures that no expert exceeds the specified capacity limit, effectively mitigating severe load imbalances and reducing latency. Note that tokens are distributed across devices under expert and data parallelism. To avoid additional communication overhead, we apply capacity constraints to tokens within each local device, similar to the constraints used during training [14]. This ensures that all devices respect the limits, maintaining strict control over token flow to the experts.
@@
where $s_{ij}$ denotes the importance score of the mapping from the $i$-th token to the $j$-th expert. With this score, each overflowed expert selectively discards those with lower scores. Let $\mathcal{J}$ be the set of overflowed experts:
$\tau_{\mathcal{J}}=\text{KthValue}(\mathcal{S}_{\mathcal{J}},C),$ (9)
where $\tau_{\mathcal{J}}$ represents the thresholds, i.e., $C$-th highest value in $\mathcal{S}_{\mathcal{J}}$, serving as a threshold to filter out excess tokens:
$T_{\mathcal{J}}\leftarrow\left\{(t,j)\;\middle|\;t\in[1,\dots,N],\;j\in\mathcal{J},\;\mathcal{S}[t,j]\geq\tau_{\mathcal{J}}[j]\right\}$ (10)
$\mathcal{S}_{\mathcal{J}}\leftarrow\mathcal{S}_{\mathcal{J}}\odot M_{\mathcal{J}},\text{where}~{}M_{\mathcal{J}}\leftarrow\mathbb{1}\left[\mathcal{S}_{\mathcal{J}}\geq\tau_{\mathcal{J}}\right],$ (11)
where $T_{\mathcal{J}}$ denotes the token indices retained by the experts indexed in $\mathcal{J}$. The scores of rejected tokens are masked to prevent them from being routed to their corresponding overflowed experts.
Regarding the specific scoring function, we explore multiple metrics and summarize them as follows:
Order: Discarding later tokens once earlier tokens have filled the expert capacity. This strategy was first introduced in Switch-Transformer [14] during training, and we extend it to the inference phase.
@@
Among these metrics, “Order” and “Reverse Order” are unstable, as shuffling sequences within a batch may result in different tokens being dropped [17]. “Random” assumes all tokens have an equal probability of being dropped. In contrast, “Score” is stable, unaffected by sequence order within a batch. Notably, there is virtually no additional computational overhead associated with calculating these metrics, and the dropping operation incurs minimal cost compared to the intensive computations performed by the experts.
(a) Token Drop
(b) Expanded Drop
Figure 3: Illustration of Capacity-Aware Token Drop (a) and Expanded Drop (b). Both methods first select experts based on gating scores. In Token Drop, tokens exceeding the local device capacity are discarded prior to All-to-All communication. Expanded Drop enhances expert utilization by allowing each token to consider additional $m$ candidate experts on the same device while still enforcing strict local capacity constraints.
Expanded Drop Enhances the Utilization of Low-load Experts
Token Drop exclusively targets overloaded experts by discarding overflowed tokens that exceed expert capacity but does not address the underutilization of low-load experts. Next, we introduce Expanded Drop to ensure a more balanced token-to-expert allocation.
A naive approach to rerouting under-selected tokens is to mask the mapping scores of overflowed experts and then reselect experts for these tokens. However, the reselection may still result in overflows, necessitating multiple rounds of selection and dropping, which increases latency. Moreover, the repeated selection and dropping substantially raise the cost of token-to-expert mapping.
Expanded Drop adopts a simple yet effective strategy: for each token, it selects additional candidate experts. Given $m$ experts deployed on a single GPU, a token not only selects the top-$k$ experts based on gating scores, but also includes $m$ local experts (e.g., 8 experts per device under 8-way expert parallelism across 8 GPUs for a total of 64 experts) for substitution if the initially selected experts are overflowed. As a result, each token may select up to $m+k$ experts. The final selection is then refined as experts drop tokens as needed to satisfy capacity constraints. This makes no change in the token assignments in experts that are overflowed by the top-$k$ experts. Meanwhile, for under-utilized experts, the expanded top-$k+m$ candidate pool increases the likelihood of receiving additional tokens. After top-$k+m$ selection and dropping, there might exists tokens that select more than $k$ experts. Through empirical analysis (Appendix D), we choose not to enforce a constraint that limits each token to selecting at most $k$ experts, thereby removing the need to explicitly retain the top-$k$ experts at the end.
Notably, the extra cost of token routing is minimal; the only difference lies in the negligible cost in the concatenation of the gating scores from either the top-$k$ or $m$ experts on the local device. Moreover, processing expanded tokens within local device eliminates inter-device communication.